### v1

# Service Rate Regions of MDS Codes & Fractional Matchings in Quasi-uniform Hypergraphs

Abstract
The service rate region (SRR) has emerged as a critical performance metric for distributed systems that store data redundantly. It measures the system’s ability to serve multiple users concurrently. Mathematically, the SRR is a polytope in $\mathbb{R}^{k}$ where each dimension corresponds to the service request rate of one of the $k$ data objects. This paper focuses on systems employing a class of Maximum Distance Separable (MDS) codes. For each code in the class, we characterize the $k$ axes intercept points of its SRR, and the smallest standard simplex that includes the SRR. We use these results to show that the SRR grows with the increasing number of systematic columns in the generator matrices. We establish a graph-theoretic framework associating this SRR problem with fractional matchings in quasi-uniform hypergraphs. Identifying the SRR polytope is equivalent to determining a particular image of the fractional-matching polytope. We introduce a notion of Greedy Matching and show that it is sufficient to focus on these matchings to characterize the SRR rather than the entire matching polytope. With these tools, we determine the SRR of a large subset of the considered class of codes. Our results generalize previous characterizations of systematic and non-systematic MDS-coded systems, offering a unified framework for analyzing service rate regions of codes.
Index Terms:
service rate, fractional matchings, recovery graph, recovery set, duality theorem
I Introduction
Distributed storage systems use erasure coding to ensure data reliability and availability. Encoded data objects are distributed over multiple storage nodes. Redundancy protects against node failures and can help mitigate download slowdowns due to slow or non-responsive nodes, known as stragglers. Traditionally, research on storage codes has focused on optimizing storage efficiency (minimizing overhead) and data recovery efficiency (reducing repair bandwidth or repair degree). See a recent monograph on the topic [1].
The service rate region (SRR) has recently emerged as a critical performance metric (see, e.g., [2] and references therein). Consider a distributed system that stores $k$ distinct data objects redundantly over $n$ servers. Each server can handle requests at a rate $\mu$. The request for object $i$ arrives at rate $\lambda_{i}$ for $i=1,\dots,k$. SRR of a storage system is the set of all data request rate vectors $(\lambda_{1},\lambda_{2},\dots,\lambda_{k})$ that the system can concurrently serve while ensuring that the total request rate allocated to any server does not exceed $\mu$. Geometrically, the SRR of this system is a convex polytope in $\mathbb{R}^{k}$, where the axes correspond to the request rates for different data objects [3].
Current research has primarily dealt with characterizing SRRs of given coding schemes. In particular, the SRR of the Simplex codes and the specific MDS were found in [2, 4], those of the Reed-Muller first-order (RM) codes in [5], and of the general RM codes in [6]. Some general characteristics of the SRR polytope are reported in [7]. The second line of inquiry is concerned with designing codes that maximize or cover a target SRR while balancing constraints such as minimal server count or code field size [8]. This paper characterizes the SRRs for a large class of MDS codes. Building upon prior research that framed SRR analysis using graph-theoretic and combinatorial methods, we significantly extend this framework to systematically investigate MDS codes with varying numbers of coded and uncoded (systematic) servers.
Distributed storage allocation problems (e.g., [9]) can often be reformulated as equivalent hypergraph matching problems (see [10] for some examples), which can then be relaxed to fractional matchings and solved efficiently via linear programming. However, these problems become challenging when the hyperedges in the graph exhibit intense and intractable overlaps or when enumerating all hyperedges is itself difficult, as noted in [6].
We here leverage the concept of recovery hypergraphs, first defined in [4], to translate the problem of analyzing service rate regions into an equivalent graph-theoretic problem. We derive explicit geometric inner and outer bounds for these regions, providing clear insights into how different storage strategies directly influence the service rate region. In particular, by analyzing these recovery hypergraphs, we prove that increasing the number of systematic servers enlarges the SRR, highlighting a fundamental trade-off between simplicity and flexibility in data service.
Moreover, we introduce a novel data request allocation strategy called Greedy Matching and establish its optimality. Utilizing this strategy, we explicitly characterize SRRs across various system configurations, generalizing and refining existing results. Our findings offer a rigorous foundation for understanding and optimizing data service capabilities in distributed storage systems, paving the way for superior system design and more efficient data access strategies.
The remainder of the paper is organized as follows. Section II first introduces the redundant storage model and formally defines service rate region (SRR) of distributed storage systems. It then reformulates the problem in graph-theoretic and linear programming terms and introduces relevant combinatorial concepts and tools. Section III derives two geometric bounds for SRRs. Section IV uses the derived bounds to establish an inclusion theorem for SRRs, proving that SRRs strictly expand as the number of systematic nodes increases. In Section V, we propose a novel Greedy Matching allocation strategy and prove its optimality. Section VI explicitly characterizes the SRRs in various configurations of the system. Finally, Section VII concludes the paper and outlines potential directions for future research.
II Problem Formulation
We first describe a redundant storage model consisting of $n$ identical nodes (servers) that collectively store multiple data objects with redundancy. We then construct a family of underlying MDS generator matrices that govern how data objects are encoded and distributed across these servers. Finally, we introduce certain concepts necessary for our analysis.
When we first mention concepts well known in the literature, we use italic. We introduce the new and less standard concepts formally through a Definition . The matrices and standard basis vectors are denoted in bold. The finite field over a prime power $q$ is denoted by $\mathbb{F}_{q}$. A linear code $\mathcal{C}$ on $\mathbb{F}_{q}$, denoted $[n,k,d]_{q}$, is a subspace of dimension $k$ with a minimum distance $d$ of the vector space of dimensions $n$ $\mathbb{F}_{q}^{n}$. The symbols $\boldsymbol{0}_{k}$ and $\boldsymbol{1}_{k}$ denote the all-zero and all-one column vectors of length $k$, respectively. The standard basis (column) vector with a 1 at position $i$ and zero elsewhere is represented by $\boldsymbol{e}_{i}$. The set of positive integers not exceeding $i$, which is $\{1,2,\dots,i\}$, is denoted $[i]$. Finally, $(x)^{+}=\max\{x,0\}$ for any real number $x$.
II-A Redundant Storage using MDS Codes
Consider a storage system that stores $k\geq 2$ data objects on $n\geq k$ servers, labeled $1,\dots,n$. We assume that all data objects have the same size and that each server has a storage capacity of one object, that is, unit storage capacity. The assumption allows us to mathematically represent objects as elements of some finite field $\mathbb{F}_{q}$. Each server stores a linear combination of the data objects in $\mathbb{F}_{q}$. The system can therefore be specified by a matrix $\boldsymbol{G}\in\mathbb{F}_{q}^{k\times n}$, called the generator matrix. If $\mathrm{o}=(o_{1},\dots,o_{k})$ is a row vector in $\mathbb{F}_{q}^{n}$ of data objects, then the $j$-th coordinate of $\mathrm{o}\cdot\boldsymbol{G}$ for $j\in\{1,2,\dots,n\}$ is the coded object stored on the $j$-th server. Each column of $\boldsymbol{G}$ corresponds to a server, and if it is equal to one of the standard basis vectors, we call the corresponding server systematic. Otherwise, we call the server coded.
We construct a family of generator matrices MDS codes over $\mathbb{F}_{q}$ with varying numbers of systematic columns. Our goal is to understand and quantify how these variations impact the service rate region.
The construction is as follows: start with a $k\times(k+n)$ MDS matrix $\boldsymbol{M}$ over a finite field $\mathbb{F}_{q}$ where $q$ is a prime or prime power such that $q\geq k+n+1$. Its structure is given by
$\boldsymbol{M}=\Bigl{[}\boldsymbol{e}_{1}\mid\boldsymbol{e}_{2}\mid\dots\mid\boldsymbol{e}_{k}\mid p_{1}\mid\dots\mid p_{n}\Bigr{]},$
where the first $k$ columns, $\boldsymbol{e}_{j}$, are standard basis vectors in $\mathbb{F}_{q}^{k}$. The remaining $n$ columns, $p_{1},\dots,p_{n}$, are parity-check columns. Because its leftmost $k$ columns form the identity matrix $\boldsymbol{I}_{k}$, $\boldsymbol{M}$ is called systematic. The construction and existence of such systematic MDS matrices over $\mathbb{F}_{q}$ have been shown, for example, in [1, 11].
From $\boldsymbol{M}$, we derive a family of $k\times n$ matrices $\boldsymbol{G}_{i}(n,k)$, indexed by $i$ where $0\leq i\leq k$. Each $\boldsymbol{G}_{i}(n,k)$ is obtained by selecting the $i$ leftmost systematic columns and the $n-i$ rightmost parity columns of $\boldsymbol{M}$:
$\boldsymbol{G}_{i}(n,k)\;=\;\Bigl{[}\boldsymbol{e}_{1}\mid\boldsymbol{e}_{2}\mid\dots\mid\boldsymbol{e}_{i}\mid p_{i+1}\mid\dots\mid p_{n}\Bigr{]}.$
We will often write $\boldsymbol{G}_{i}$ instead of $\boldsymbol{G}_{i}(n,k)$, and denote by $G_{i}^{l}$ (without boldface) the $l$-th column of $\boldsymbol{G}_{i}$. Since $\boldsymbol{M}$ is an MDS matrix, any $k$ columns of $\boldsymbol{M}$ are linearly independent, and the same holds for the $k$ columns of $\boldsymbol{G}_{i}$. Hence each $\boldsymbol{G}_{i}$ also generates an MDS code. Moreover, $\boldsymbol{G}_{i}$ and $\boldsymbol{G}_{i+1}$ differ in exactly one column, namely, the $(i+1)$-th column. These observations will play a key role in proving the inclusion relations for service regions (see Section IV).
We use matrices $\boldsymbol{G}_{i}$ as generator matrices for our system. Practically, in a storage system generated by $\boldsymbol{G}_{i}$, each of the first $i$ servers stores the uncoded copy of a single data object (raw data). The remaining $n-i$ servers store linearly coded copies.
II-B Recovery Sets and Service Rate Region
A recovery set for object $o_{i}$ is a set of stored symbols that can be used to recover $o_{i}$. A subset $R$ of columns in $\boldsymbol{G}$ is a recovery set for data object $o_{j}$ (aka recovery set for vector $\boldsymbol{e}_{j}$), if
$\boldsymbol{e}_{j}\in\textsf{span}(\cup_{i\in R}\{\boldsymbol{c}_{i}\})\triangleq\textsf{span}(R),\,\text{ and }\,\boldsymbol{e}_{j}\notin\textsf{span}(S)\text{ for any proper subset }S\subsetneq R.$
In other words, $R$ is a minimal set of columns whose span in $\mathbb{F}_{q}$ includes $\boldsymbol{e}_{j}$.
Let $\mathcal{R}_{i}=\{R_{i,1},\ldots,R_{i,t_{i}}\}$ be the $t_{i}$ recovery sets for the object $o_{i}$. Define $\mu_{l}\in\mathbb{R}_{\geq 0}$ as the average rate at which the server $l\in[n]$ processes requests for data objects, also referred to as its capacity. The vector $\boldsymbol{\mu}=(\mu_{1},\ldots,\mu_{n})$ represents the service capacities of all servers. Furthermore, requests for an object $o_{i}$ arrive at a rate $\lambda_{i}$ for all $i\in[k]$, with the request rates for all objects represented by the vector $\boldsymbol{\lambda}=(\lambda_{1},\ldots,\lambda_{k})\in\mathbb{R}_{\geq 0}^{k}$.
Consider the class of scheduling strategies that assign a fraction of requests for an object to each of its recovery sets. Let $\lambda_{i,j}$ be the portion of requests for object $o_{i}$ that are assigned to the recovery set $R_{i,j},j\in[t_{i}]$. The service rate region (SRR) $\mathcal{S}(\boldsymbol{G})\subset\mathbb{R}_{\geq 0}^{k}$ is defined as the set of all request vectors $\boldsymbol{\lambda}$ that can be served by a coded storage system with generator matrix $\boldsymbol{G}$ and service rate $\boldsymbol{\mu}$. Such vectors $\boldsymbol{\lambda}$ are called achievable. Therefore, $\mathcal{S}(\boldsymbol{G})$ is the set of all vectors $\boldsymbol{\lambda}$ for which there exist $\lambda_{i,j}\in\mathbb{R}_{\geq 0},i\in[k]$ and $j\in[t_{i}]$, satisfying the following constraints:
$\sum_{j=1}^{t_{i}}\lambda_{i,j}$ $=\lambda_{i},\quad\forall\,i\in[k],$ (1)
$\sum_{i=1}^{k}\sum_{\begin{subarray}{c}j=1\\ l\in R_{i,j}\end{subarray}}^{t_{i}}\lambda_{i,j}$ $\leq\mu_{l},\quad\forall\,l\in[n],$ (2)
$\lambda_{i,j}$ $\in\mathbb{R}_{\geq 0},\quad\forall\,i\in[k],j\in[t_{i}].$ (3)
The constraints (1) guarantee that the demands for all objects are satisfied, and constraints (2) ensure that no server receives requests at a rate larger than its service capacity. Such vectors $\boldsymbol{\lambda}$ form the service polytope in $\mathbb{R}_{\geq 0}^{k}$. For any vector $\boldsymbol{\lambda}$, a set $\{\lambda_{i,j}:1\leq i\leq k,1\leq j\leq t_{i}\}$ that satisfies (1)–(3) is referred to as a valid allocation associated with $\boldsymbol{\lambda}$.
In $\boldsymbol{G}_{i}(n,k)$, recovery sets for each basis vector $\boldsymbol{e}_{j}$ follow: (i) If $i<j$, any $k$-subset of columns forms a non-systematic recovery set for $\boldsymbol{e}_{j}$. (ii) If $i\geq j$, the $j$-th column equals $\boldsymbol{e}_{j}$, giving a systematic recovery set of size $1$, while any $k$-subset not containing the $j$-th column serves as a non-systematic recovery set.
Remark 1.
All non-systematic recovery sets have size $k$, since no smaller set excluding $\boldsymbol{e}_{j}$ can span it without violating the MDS property. Hence, the size of a recovery set is either $1$ or $k$.
II-C Recovery Hypergraphs
A hypergraph (or simply graph) is a pair $(V,E)$, where $V$ is a finite set of vertices, and $E$ is a multiset whose elements are subsets of $V$, called hyperedges (or edges). The size of an edge is its cardinality. A hypergraph is $k$-uniform if each edge has size exactly $k$. We further generalize uniformity to define a $(k,r)$-quasi-uniform hypergraph:
Definition 1.
A hypergraph is called $(k,r)$-quasi-uniform if each hyperedge has size either $k$ or $r$.
We next introduce the concept of recovery hypergraph, first defined in [4], associated with each generator matrix. This notion enables us to reformulate our problem in terms of graph theory, allowing us to leverage its well-developed tools and analytical results.
For the matrix $\boldsymbol{G}_{i}(n,k)$, we define its recovery hypergraph $\Gamma_{i}(n,k)$ (vertices and edges) as follows. $\Gamma_{i}(n,k)$ has $n+i$ vertices:
  • $n$ vertices, each corresponding to one distinct column of $\boldsymbol{G}_{i}(n,k)$,
  • $i$ additional vertices $(\boldsymbol{0}_{k})_{j}=[\begin{smallmatrix}0,\dots,0\end{smallmatrix}]^{T}_{j}$, for $j=1,\dots,i$, each representing a length-$k$ zero vector.
Next, we precisely describe how vertices form hyperedges in the recovery hypergraph. A set of vertices forms a hyperedge labeled $\boldsymbol{e}_{j}$ if their corresponding columns constitute a recovery set for the basis vector $\boldsymbol{e}_{j}$. Specifically, if a systematic column $\boldsymbol{e}_{j}$ appears in the matrix $\boldsymbol{G}_{i}$, the vertex corresponding to this systematic column is called a systematic vertex. Each systematic vertex is directly connected to an additional zero vertex $\bigl{[}\begin{smallmatrix}0,\dots,0\end{smallmatrix}\bigr{]}^{T}_{j}$, forming an edge labeled $\boldsymbol{e}_{j}$, referred to as a systematic edge. Edges formed by any minimal recovery set that does not include a systematic vertex are called non-systematic edges. Thus, every edge is either systematic (of size 2) or non-systematic (of size $k$). Additionally, the same vertex set may form multiple parallel edges with distinct labels if it can serve as recovery sets for multiple basis vectors. However, each hyperedge carries exactly one unique label.
Figure 1 illustrates the matrices $\boldsymbol{G}_{i}(4,2)$ along with their corresponding recovery hypergraphs $\Gamma_{i}(4,2)$ for $i=0,1,2$. In these matrices, the four parity columns are explicitly given by:
$\bigl{[}\begin{smallmatrix}1\\ 1\end{smallmatrix}\bigr{]},\bigl{[}\begin{smallmatrix}1\\ \alpha\end{smallmatrix}\bigr{]},\bigl{[}\begin{smallmatrix}1\\ \alpha^{2}\end{smallmatrix}\bigr{]},\bigl{[}\begin{smallmatrix}1\\ \alpha^{3}\end{smallmatrix}\bigr{]}.$
In $\boldsymbol{G}_{0}(4,2)$, for example, vertices $\bigl{[}\begin{smallmatrix}1\\ 1\end{smallmatrix}\bigr{]}$ and $\bigl{[}\begin{smallmatrix}1\\ \alpha\end{smallmatrix}\bigr{]}$ form two parallel edges, one labeled $\boldsymbol{e}_{1}$ and the other $\boldsymbol{e}_{2}$.
Figure 1: $\boldsymbol{G}_{i}(4,2)$ matrices with their recovery hypergraphs $\Gamma_{i}(4,2)$, $i=0,1,2$.
Since $\boldsymbol{G}_{0}$ is an MDS matrix without systematic columns, its recovery hypergraph $\Gamma_{0}(n,k)$ is a $k$-uniform hypergraph composed exclusively of non-systematic edges. Additionally, by Lemma 1, for $i=1,\dots,k$, each systematic edge in $\Gamma_{i}(n,k)$ has size 2, while all non-systematic edges have size $k$. Hence, $\Gamma_{i}(n,k)$ forms a $(k,2)$-quasi-uniform hypergraph. Moreover, as every set of $k$ columns is linearly independent and capable of recovering any basis vector $\boldsymbol{e}_{j}$, the vertices representing these columns form $k$ parallel non-systematic edges, each uniquely labeled by distinct $\boldsymbol{e}_{j}$. For any hypergraph $\Gamma$ and subset $I\subseteq[k]$, the $I$-induced subgraph of $\Gamma$ includes exactly those edges labeled by basis vectors $\boldsymbol{e}_{j}$ with $j\in I$, together with all vertices incident to these edges.
II-D Fractional Matching and Service Polytopes for Recovery Graphs
A fractional matching in a hypergraph $(V,E)$ is a vector $\boldsymbol{w}\in\mathbb{R}_{\geq 0}^{|E|}$ whose components $w_{\epsilon}$, for $\epsilon\in E$, are nonnegative and satisfy
$\sum_{\epsilon\,\ni\,v}w_{\epsilon}\;\leq\;1\quad\text{for each vertex }v\in V.$
The set of all fractional matchings in $\Gamma_{G}=(V,E)$ forms a polytope in $\mathbb{R}_{\geq 0}^{|E|}$, called the fractional matching polytope, denoted $\mathrm{FMP}(\Gamma_{G})$. It can be written as
[…]
where $\boldsymbol{A}$ is the $|V|\times|E|$ incidence matrix of $\Gamma_{G}$, $\boldsymbol{1}$ is the all-one vector of length $|V|$, and $\boldsymbol{0}$ is the all-zero vector of length $|E|$.
Figure 2 illustrates two distinct matchings that result in the same service vector $\boldsymbol{\lambda}=(1.5;0.75)$. That is, for these matchings, the sum of weights $w_{\epsilon}$ over all edges $\epsilon$ labeled $\boldsymbol{e}_{j}$ equals $\lambda_{j}$ for $j=1,2$. These two matchings embody two different valid allocations for a single vector $\boldsymbol{\lambda}$.
By definition, each matching ensures that the total request assigned to each node does not exceed its capacity $\mu=1$. Therefore, $\boldsymbol{\lambda}\in\mathcal{S}(\boldsymbol{G})$. A request vector $\boldsymbol{\lambda}$ lies within the SRR of a storage system employing a code $\boldsymbol{G}$ if and only if there exists a matching $\boldsymbol{w}$ such that the sum of weights on edges labeled $\boldsymbol{e}_{j}$ equals $\lambda_{j}$, as established in the following result.
Definition 2.
Consider a system employing an $[n,k]$ code with generator matrix $\boldsymbol{G}$ and uniform server availability, that is, $\boldsymbol{\mu}=\boldsymbol{1}_{n}$. The service rate for $\boldsymbol{e}_{j}$ under a fractional matching $\boldsymbol{w}$, denoted $\lambda_{j}(\boldsymbol{w})$, is the sum of the weights $w_{\epsilon}$ over all hyperedges $\epsilon$ labeled by $\boldsymbol{e}_{j}$. The corresponding service vector is given by
$\boldsymbol{\lambda}(\boldsymbol{w})=(\lambda_{1}(\boldsymbol{w}),\dots,\lambda_{k}(\boldsymbol{w})).$
Each $\boldsymbol{w}$ defines a valid allocation for $\boldsymbol{\lambda}$.
Proposition 1 ([2], Proposition 1).
$\boldsymbol{\lambda}=(\lambda_{1},\dots,\lambda_{k})$ is achievable if and only if there exists a fractional matching $\boldsymbol{w}$ in the recovery graph $\Gamma_{G}$ such that
$\boldsymbol{\lambda}=\boldsymbol{\lambda}(\boldsymbol{w}).$
[…]
Proposition 1 shows that $\mathcal{S}(\boldsymbol{G})$ is the image of $\mathrm{FMP}(\Gamma_{G})$ under a linear map from $\mathbb{R}^{|E|}$ to $\mathbb{R}^{k}$. Specifically, let $\boldsymbol{S}$ be the $|E|\times k$ matrix whose entries are
$[\boldsymbol{S}]_{\epsilon,j}\;=\;\begin{cases}1,&\text{if edge $\epsilon$ is labeled by $\boldsymbol{e}_{j}$,}\\[6.0pt] 0,&\text{otherwise}.\end{cases}$
Since each edge is associated with exactly one label, each row of $\boldsymbol{S}$ contains exactly one entry equal to 1, with all other entries equal to 0. For a matching $\boldsymbol{w}\in\mathrm{FMP}(\Gamma_{G})$, the resulting service vector is $\boldsymbol{\lambda}(\boldsymbol{w})=\boldsymbol{w}\,\boldsymbol{S}$. If $\boldsymbol{\lambda}\in\mathcal{S}(\boldsymbol{G})$, there may be multiple $\boldsymbol{w}\in\mathrm{FMP}(\Gamma_{G})$ such that $\boldsymbol{\lambda}=\boldsymbol{w}\,\boldsymbol{S}$.
The service rate region $\mathcal{S}(\boldsymbol{G})$, therefore, also forms a polytope in $\mathbb{R}_{\geq 0}^{k}$, and we use the terms service polytope and service rate region interchangeably. We denote the SRR corresponding to $\boldsymbol{G}_{i}(n,k)$ by $\mathcal{S}_{i}(n,k)$ or simply $\mathcal{S}_{i}$.
Remark 2.
From the above argument, Proposition 1 can also be stated as follows:
[…]
$\boldsymbol{\lambda}$ $=\boldsymbol{w}\boldsymbol{S},$ (4)
[…]
$\boldsymbol{w}$ $\geq\boldsymbol{0}_{n}.$
When $\boldsymbol{\lambda}$ is achievable, the problem might have infinitely many solutions. Such solutions lie in a space whose dimension is $|E|-\mathrm{rank}(\boldsymbol{S})\geq|E|-k$.
The size of a matching $\boldsymbol{w}$ is $\sum\limits_{\epsilon\in E}w_{\epsilon}$. The matching number $\nu^{*}(\Gamma_{G})$ is the maximum possible matching size:
$\nu^{*}(\Gamma_{G})\;=\;\max_{\boldsymbol{w}\in\mathrm{FMP}(\Gamma_{G})}\sum_{\epsilon\,\in\,E}w_{\epsilon}.$
A fractional vertex cover of $(V,E)$ is a vector $\boldsymbol{w}\in\mathbb{R}^{|V|}$ with nonnegative components $w_{v}$ such that $\sum_{v\in\epsilon}w_{v}\geq 1$ for every edge $\epsilon\in E$. Its size is $\sum\limits_{v\in V}w_{v}$. The vertex cover number $\tau^{*}(\Gamma_{G})$ is the minimum size of any fractional vertex cover:
[…]
Finding $\nu^{*}(\Gamma_{G})$ is a linear program whose dual problem finds the minimum fractional vertex cover $\tau^{*}(\Gamma_{G})$. By the strong Duality theorem, $\nu^{*}(\Gamma_{G})\;=\;\tau^{*}(\Gamma_{G}).$
Framing the SRR problem within graph theory allows us to leverage known results from the literature. The former is seen in Proposition 1, while the latter follows from the next result, which can easily be proved using Proposition 1.
Proposition 2.
For any vector $\boldsymbol{\lambda}=(\lambda_{1},\lambda_{2},\dots,\lambda_{k})$ in the service region $\mathcal{S}(\boldsymbol{G})$,
$\tau^{*}(\Gamma_{G})\;\geq\;\sum_{i=1}^{k}\lambda_{i}.$ (5)
Proof.
Since $\boldsymbol{\lambda}\in\mathcal{S}(\boldsymbol{G})$, by Proposition 1 there exists a matching $\boldsymbol{w}$ such that $\boldsymbol{\lambda}=\boldsymbol{\lambda}(\boldsymbol{w})=\boldsymbol{w}\boldsymbol{S}$. Denote by $\boldsymbol{1}_{k}$ the length-$k$ column vectors of all ones. Then,
$\sum_{j=1}^{k}\lambda_{j}(\boldsymbol{w})\;=\;\boldsymbol{\lambda}(\boldsymbol{w})\cdot\boldsymbol{1}_{k}\;=\;\boldsymbol{w}(\boldsymbol{S}\cdot\boldsymbol{1}_{k})\;=\;\boldsymbol{w}\cdot\boldsymbol{1}_{k}\;=\;\sum_{\epsilon\in E}w_{\epsilon}\;\leq\;\max_{\boldsymbol{w}\in\mathrm{FMP}(\Gamma_{G})}\sum_{\epsilon\in E}w_{\epsilon}\;=\;\nu^{*}(\Gamma_{G})\;=\;\tau^{*}(\Gamma_{G}).$
(The third equality follows from the fact that each row of $\boldsymbol{S}$ has exactly one entry equal to 1, with all others equal to 0. The second-to-last equality follows from the definition of $\nu^{*}(\Gamma_{G})$, and the last equality comes from the LP duality.) ∎
Remark 3.
This result generalizes Proposition 2 in [2], where the authors proved that the size of any integral vertex cover provides an upper bound on the achievable heterogeneous sum rate. It establishes a fundamental bound on the sum of request rates within the service polytope of any linearly coded storage system.
Consequently, the size of any (fractional) vertex cover serves as an upper bound on the sum rate for any achievable vector $\boldsymbol{\lambda}$. In particular, for large hypergraphs with specific structures, such as when the number of vertices $n$ is significantly smaller than the number of edges $\binom{n}{k}$ or when most edges have a large cardinality, this lemma provides a simple, yet tight estimate of the achievable sum rate.
Remark 4.
A straightforward manipulation of Proposition 2 shows that it also applies to any subgraph of $\Gamma_{G}$. Specifically, if $I$ is any subset of $[k]$ and $\Gamma^{\prime}$ is the $I$-induced subgraph of $\Gamma$, then
$\tau^{*}(\Gamma^{\prime})\;\geq\;\sum_{i\in I}\lambda_{i}.$
Main Notation Summary
  • $k$
    number of data objects.
  • $n$
    number of servers (storage nodes).
  • $i$
    number of systematic columns (or systematic nodes) in the MDS generator matrix.
  • $\lambda_{j}$
    the rate of requests for the $j$-th object.
  • $\boldsymbol{\lambda}$
    vector of request rates $\boldsymbol{\lambda}=(\lambda_{1},\lambda_{2},\dots,\lambda_{n})$
  • $\boldsymbol{G}$
    Generator matrix.
  • $\mathcal{S}(\boldsymbol{G})$
    Service rate region (SRR) of the system using code $\boldsymbol{G}$.
  • $R$
    A recovery set for a data object (basis vector) in the code.
  • $\Gamma_{G}$
    Recovery hypergraph associated with the generator matrix $\boldsymbol{G}$.
  • w
    Fractional matching weight vector.
  • $\nu^{*}(\Gamma_{G})$
    Fractional matching number the hypergraph $\Gamma_{G}$.
  • $\tau^{*}(\Gamma_{G})$
    Fractional vertex cover number of the hypergraph $\Gamma_{G}$. We have $\nu^{*}(\Gamma_{G})=\tau^{*}(\Gamma_{G})$.
  • +$\mathbb{R}^{k}_{+}$
    positive orthant of $\mathbb{R}^{k}$.
III Two Bounding Simplices
The simplices of interest in $\mathbb{R}^{k}$ are convex hulls of the origin and the $k$ axis-aligned points $c_{j}\boldsymbol{e}_{j}$, for $j=1,\dots,k$, where each $c_{j}>0$. When all $c_{j}$ are equal to a common constant $c\geq 0$, the resulting simplex corresponds to the set of points $(x_{1},x_{2},\dots,x_{k})\in\mathbb{R}^{k}_{+}$ satisfying $\sum_{j=1}^{k}x_{j}\leq c.$
This section introduces two simplex polytopes related to any SRR: the Maximal matching simplex and the Maximal achievable simplex. The former is the smallest simplex that contains the SRR, while the latter is the largest simplex fully contained within it. These simplices help localize the SRR, in general. In Sec. IV, we will use them to derive an inclusion theorem for $\mathcal{S}_{i}(n,k)$ that shows the role of systematic nodes.
III-A Maximal Matching Simplex
For each $\mathcal{S}_{i}(n,k)$, we seek an outer simplex containing the service rate region (SRR), defined as
$\sum_{j=1}^{k}\lambda_{j}\;\leq\;c,~{}~{}\lambda_{j}\geq 0.$
where $c$ is a constant specific to $\mathcal{S}_{i}(n,k)$. Proposition 2 provides this bound. Recall from Eq. (5) that for any generator matrix $\boldsymbol{G}_{i}(n,k)$ and any service vector $\boldsymbol{\lambda}\in\mathcal{S}_{i}(n,k)$, the SRR lies within the positive orthant simplex defined by
$\sum_{j=1}^{k}\lambda_{j}\;\leq\;\tau^{*}(\Gamma_{i}(n,k)=\nu^{*}(\Gamma_{i}(n,k)).$
We refer to this region as the maximal matching simplex. We next determine the value of $\tau^{*}(\Gamma_{i}(n,k))$ by constructing explicit matchings and vertex covers in two distinct cases.
Case 1: $n-i\geq k$
Consider the following vertex cover of $\Gamma_{i}(n,k)$:
$w_{v}\;=\;\begin{cases}1,&\text{if $v$ is a systematic vertex},\\[4.0pt] {1}/{k},&\text{otherwise}.\end{cases}$
We show that this is a valid vertex cover.
  • For a systematic edge $\epsilon$, there is exactly one vertex $v$ with $w_{v}=1$.
  • For a non-systematic edge $\epsilon$, Lemma 1 guarantees that $\epsilon$ has size $k$. Hence, $\sum\limits_{v\in\epsilon}w_{v}=k\cdot\tfrac{1}{k}=1$.
Thus, the total weight assigned to the vertices incident to each edge is at least 1, so this indeed forms a vertex cover of size $i+\dfrac{n-i}{k}$.
On the other hand, we construct a matching by assigning:
$w_{\epsilon}\;=\;\begin{cases}1,&\text{if edge $\epsilon$ is systematic},\\[3.0pt] 0,&\text{if $\epsilon$ is non-systematic and contains a systematic vertex},\\[3.0pt] \bigl{(}k\,\binom{n-i-1}{k-1}\bigr{)}^{-1},&\text{otherwise}.\end{cases}$
We check that this matching is valid as follows
  • A systematic vertex $v$ is incident to exactly one systematic edge with weight 1.
  • A non-systematic vertex $v$ belongs to $\binom{n-i-1}{k-1}$ different size-$k$ sets of columns that do not include any systematic column. Each such set corresponds to $k$ parallel edges, each assigned weight $\bigl{(}k\,\binom{n-i-1}{k-1}\bigr{)}^{-1}$. Summing over all these edges gives
[…]
Hence, the total weight assigned to the edges incident to each vertex $v$ does not exceed 1.
Counting the edges:
  • There are $i$ systematic vertices, each is incident with one systematic edge.
  • Excluding these, the remaining $n-i$ non-systematic vertices form a $k$-uniform hypergraph. Because $n-i\geq k$, there are $k\,\binom{n-i}{k}$ edges of size $k$ in this graph.
The total matching size is
$i\;+\;\left(k\,\binom{n-i-1}{k-1}\right)^{-1}\,\cdot\,\left(k\,\binom{n-i}{k}\right)\;=\;i+\frac{n-i}{k}.$
By LP duality, since there is a vertex cover and a matching with the same size,
$\nu^{*}(\Gamma_{i}(n,k))\;=\;\tau^{*}(\Gamma_{i}(n,k))\;=\;i+\frac{n-i}{k}.$ (6)
Thus, in this case, the Maximal matching simplex is
$\begin{cases}\boldsymbol{0}\leq\boldsymbol{\lambda},\\[5.0pt] \sum_{j=1}^{k}\lambda_{j}\;\leq\;i+\dfrac{n-i}{k}.\end{cases}$ (7)
Note that when $i=0$, $\Gamma_{0}(n,k)$ reduces to the $k$-uniform hypergraph of $n$ vertices, where
$\nu^{*}(\Gamma_{0}(n,k))=\tau^{*}(\Gamma_{0}(n,k))={n}/{k}.$ (8)
Case 2: $n-i<k$
In this regime, consider the following vertex cover in $\Gamma_{i}(n,k)$:
$w_{v}\;=\;\begin{cases}1,&\text{if $v$ is systematic},\\ 0,&\text{otherwise}.\end{cases}$
Because $n-i<k$, any non-systematic edge has size $k$ and must include at least one systematic vertex, so all edges are covered. The total weight is $i$. Next, define a matching:
$w_{\epsilon}\;=\;\begin{cases}1,&\text{if $\epsilon$ is systematic},\\ 0,&\text{otherwise}.\end{cases}$
Since the systematic edges do not overlap on any vertex, they form a valid matching of size $i$. Therefore, by duality,
$\nu^{*}(\Gamma_{i}(n,k))\;=\;\tau^{*}(\Gamma_{i}(n,k))\;=\;i.$ (9)
Hence, in this case, the Maximal matching simplex is
$\begin{cases}\boldsymbol{0}\leq\boldsymbol{\lambda},\\[3.0pt] \sum_{j=1}^{k}\lambda_{j}\;\leq\;i.\end{cases}$ (10)
III-B Axes-Intercept Points and the Maximal Achievable Simplex
We now characterize an inner simplex contained in $\mathcal{S}(\boldsymbol{G})$. The axes-intercept vertices of this simplex coincide with those of $\mathcal{S}(\boldsymbol{G})$. For each $j\in[k]$, define
$\lambda_{j}^{\text{int}}\triangleq\max_{\gamma\cdot\boldsymbol{e}_{j}\in\mathcal{S}(\boldsymbol{G})}\gamma.$
That is, $\lambda_{j}^{\text{int}}$ is the maximum achievable demand for the object $j$ when all other demands are zero. For example, when $j=1$, we have $\lambda_{1}^{\text{int}}=\max\{\,\gamma\mid(\gamma,0,\dots,0)\in\mathcal{S}(\boldsymbol{G})\}$.
We claim that $\lambda_{j}^{\text{int}}=\max\limits_{\boldsymbol{\lambda}\in\mathcal{S}(\boldsymbol{G})}\lambda_{j}\triangleq\lambda_{j}^{\max}$. To see that, consider $j=1$, wlog. Suppose, for contradiction, that there exists $\boldsymbol{\eta}=(\eta_{1},\eta_{2},\dots,\eta_{k})\in\mathcal{S}(\boldsymbol{G})$ with $\eta_{1}>\lambda_{1}^{\text{int}}$. Consider the vector $\boldsymbol{\eta}^{\prime}=(\eta_{1},0,\dots,0)$. Since $\boldsymbol{\eta}$ satisfies constraints (1)–(3), and these constraints remain valid when all but one coordinate are set to zero, it follows that $\boldsymbol{\eta}^{\prime}\in\mathcal{S}(\boldsymbol{G})$, contradicting the definition of $\lambda_{1}^{\text{int}}$. Therefore, $\lambda_{1}^{\text{int}}=\lambda_{1}^{\max}$, and similarly $\lambda_{j}^{\text{int}}=\lambda_{j}^{\max}$ for all $j\in[k]$.
Geometrically, $\lambda_{j}^{\text{max}}$ represents the intercept of the service polytope with the axis defined by $\boldsymbol{e}_{j}$. Practically, $\lambda_{j}^{\text{max}}$ quantifies the maximal demand $\lambda_{j}$ that our system can support. We therefore call it the maximum achievable demand for $\lambda_{j}$.
Define the simplex $\mathcal{A}$ as:
$\mathcal{A}=\textsf{conv}\Bigl{(}\bigl{\{}\boldsymbol{0}_{k},\lambda_{1}^{\text{max}}\boldsymbol{e}_{1},\lambda_{2}^{\text{max}}\boldsymbol{e}_{2},\dots,\lambda_{k}^{\text{max}}\boldsymbol{e}_{k}\bigr{\}}\Bigr{)},$
where $\textsf{conv}(\mathcal{A})$ denotes the convex hull of the set $\mathcal{A}$, defined as $\mathcal{A}=\{\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{p}\}\subset\mathbb{R}^{k}$. Specifically, $\textsf{conv}(\mathcal{A})$ consists of all convex combinations of the elements in $\mathcal{A}$, i.e., all vectors of the form
$\sum_{i=1}^{p}\gamma_{i}\boldsymbol{v}_{i},\quad\text{where }\gamma_{i}\geq 0\text{ and }\sum_{i=1}^{p}\gamma_{i}=1,$
as described in [12].
In [5, Lemma 1] it was shown that $\mathcal{S}(\boldsymbol{G})$ is a non-empty, convex, closed, and bounded subset of $\mathbb{R}_{\geq 0}^{k}$. Thus $\mathcal{A}\subseteq\mathcal{S}(\boldsymbol{G})$. This implies that all points within the simplex $\mathcal{A}$ are achievable, and we refer to it as the Maximal achievable simplex. Therefore, characterizing these extreme points is of significant interest. The following theorem helps us to achieve this.
Proposition 3.
For $\mathcal{S}_{i}(n,k)$,
$\lambda_{j}^{\text{max}}\;=\;\begin{cases}1+\dfrac{n-1}{k},&\text{if }j\leq i\text{ and }n-1\geq k,\\[6.0pt] 1,&\text{if }j\leq i\text{ and }n=k,\\[6.0pt] \dfrac{n}{k},&\text{if }j>i.\end{cases}$
Proof.
Consider the axis-intercept vector $\lambda_{j}^{\text{max}}\boldsymbol{e}_{j}$, which has all components equal to zero except $\lambda_{j}$. Since all other rates are zero, in $\Gamma_{i}(n,k)$, we remove all edges with labels different from $\boldsymbol{e}_{j}$. Let $\Gamma_{i}^{j}(n,k)$ be the subgraph of $\Gamma_{i}(n,k)$ that contains only edges labeled $\boldsymbol{e}_{j}$. Then by Remark 4,
$\lambda_{j}^{\text{max}}\;=\;\nu^{*}\bigl{(}\Gamma_{i}^{j}(n,k)\bigr{)}.$
Case 1: $j>i$. All edges labeled $\boldsymbol{e}_{j}$ are non-systematic, so $\Gamma_{i}^{j}(n,k)$ is a $k$-uniform hypergraph on $n$ vertices. By Eq. (8), $\lambda_{j}^{\text{max}}\;=\;{n}/{k}.$
Case 2: $j\leq i$. For $\boldsymbol{e}_{j}$, the subgraph $\Gamma_{i}^{j}(n,k)$ contains:
  • One systematic edge of size 1 (the column $\boldsymbol{e}_{j}$ itself).
  • Non-systematic edges of size $k$ formed by choosing any $k$-subset of columns excluding $\boldsymbol{e}_{j}$.
Hence, $\Gamma_{i}^{j}(n,k)$ splits into two disjoint hypergraphs: a 2-uniform hypergraph with 2 nodes (the systematic column and its associated zero vector) and a $k$-uniform hypergraph with $(n-1)$ nodes (all remaining columns). Therefore by Equations (6) and (9),
$\nu^{*}(\Gamma_{i}^{j})\;=\;\begin{cases}1,&\text{if }n=k,\\ 1+(n-1)/k,&\text{if }n-1\geq k,\end{cases}$
matching the desired piecewise form. ∎
Figure 3 (a) illustrates the service region $\mathcal{S}_{2}(4,2)$ along with its two bounding simplices. Some points on the Maximal Matching Simplex lie within $\mathcal{S}_{2}(4,2)$, while others lie outside, reflecting that the true service region is strictly between these two simplices.
From the derived expressions of these two simplices, we see that the Maximal Matching Simplex is always within a factor of $k$ of the Maximal Achievable Simplex. The worst case (when the two are furthest apart) occurs with $\mathcal{S}_{k}(k,k)$ (systematic encoding of $k$ data objects into $k$ servers), where the Maximal Matching Simplex is given by $\sum_{j=1}^{k}\lambda_{j}\leq k$ and the Maximal Achievable Simplex is given by $\sum_{j=1}^{k}\lambda_{j}\leq 1$, which is $k$-times smaller. Figure 3 (b) depicts this scenario with $\mathcal{S}_{2}(2,2)$.
$\lambda_{b}$$\lambda_{a}$2.52.533maximal achievable simplexmaximal matching simplexservice polytope
$\lambda_{b}$$\lambda_{a}$1122maximal achievable simplexmaximal matching simplexservice polytope
Figure 3: (Left) The service region $\mathcal{S}_{2}(4,2)$ (in red) together with its Maximal matching and Maximal achievable simplices. Some points on the Maximal matching simplex are indeed achievable. (Right) $\mathcal{S}_{2}(2,2)$ and its two bounding simplices. In this case, they are “further apart” from each other than in case (a).
We see that for MDS codes, the two bounding simplices of the SRR can differ by a factor proportional to the number of objects $k$. In contrast, for Reed–Muller codes, the bounds differ by a constant factor of at most 2, independent of the code parameters [6].
When $i=0$, $\Gamma_{0}(n,k)$ is a $k$-uniform hypergraph, so these two simplices coincide exactly with the SRR polytope. Concretely, (cf. [2])
$\mathcal{S}_{0}(n,k)=\begin{cases}\boldsymbol{0}\leq\boldsymbol{\lambda},\\ \sum_{j=1}^{k}\lambda_{j}\;\leq\;{n}/{k}.\end{cases}$
Moreover, from (7), (10) and Proposition 3, we observe that for fixed $n$ and $k$, decreasing $i$ reduces the gap between the two bounding simplices, enabling a more precise characterization of the service polytope. Similarly, for fixed $i$ and $k$, increasing $n$ produces the same effect, improving the accuracy with which the SRR can be located.
IV An Inclusion Theorem for $\mathcal{S}_{i}(n,k)$, $0\leq i\leq k$
In this section, we prove that the service rate regions of MDS codes grow strictly larger as the number of systematic columns increases. Formally:
Theorem 1.
For each $k\geq 2$ and $n\geq k$
$\mathcal{S}_{0}\;\subsetneq\;\mathcal{S}_{1}\;\subsetneq\;\mathcal{S}_{2}\;\subsetneq\;\dots\;\subsetneq\;\mathcal{S}_{k-1}\;\subsetneq\;\mathcal{S}_{k}.$ (11)
In other words, having more systematic columns strictly enlarges the SRR polytopes.
Proof.
From the definition of MDS matrices, any set of $k$ columns excluding $e_{j}$ in $\boldsymbol{G}_{i}(n,k)$ can serve as a recovery set for each basis vector $\boldsymbol{e}_{j}$. In the corresponding recovery hypergraph $\Gamma_{i}(n,k)$, the number of hyperedges labeled with $\boldsymbol{e}_{j}$ is:
$\begin{cases}1+\binom{n-1}{k},&\text{if $\boldsymbol{G}_{i}(n,k)$ contains thesystematic column $\boldsymbol{e}_{j}$, i.e.\ $i\geq j$},\\ \binom{n}{k},&\text{if $\boldsymbol{G}_{i}(n,k)$ does not contain $\boldsymbol{e}_{j}$, i.e.\ $i<j$}.\end{cases}$
Additionally, in each recovery hypergraph $\Gamma_{G}$, any systematic recovery set is dedicated to exactly one basis vector, whereas $k$-columns recovery sets of can recover any basis vector.
Comparing $\boldsymbol{G}_{i}$ and $\boldsymbol{G}_{i+1}$. By construction in Section II, $\boldsymbol{G}_{i}$ and $\boldsymbol{G}_{i+1}$ differ in exactly one column: $\boldsymbol{G}_{i+1}$ is obtained by replacing the $(i+1)$-th column in $\boldsymbol{G}_{i}$ (i.e., parity column $p_{i+1}$) with the systematic column $\boldsymbol{e}_{i+1}$. All other columns remain the same, and any systematic column in $\boldsymbol{G}_{i}$ still appears in $\boldsymbol{G}_{i+1}$.
Case Analysis on $\boldsymbol{e}_{j}$.
  • Case 1: $j>i$. Then there is no systematic recovery set for $\boldsymbol{e}_{j}$ in $\boldsymbol{G}_{i}$. There might be a systematic recovery set for $\boldsymbol{e}_{j}$ in $\boldsymbol{G}_{i+1}$, specifically if $j=i+1$. In $\boldsymbol{G}_{i}$, recovery for $\boldsymbol{e}_{j}$ is possible only via non-systematic edges of size $k$. Whenever such a recovery set contains the replaced parity column $p_{i+1}$, we can form a corresponding recovery set in $\boldsymbol{G}_{i+1}$ by substituting $\boldsymbol{e}_{i+1}$ for $p_{i+1}$. Because $\boldsymbol{G}_{i+1}$ is an MDS matrix, these new $k$-column sets are linearly independent and also recover $\boldsymbol{e}_{j}$. Hence, any service rate achievable in $\mathcal{S}_{i}$ is still feasible in $\mathcal{S}_{i+1}$.
  • Case 2: $j\leq i$. In this scenario, $\boldsymbol{e}_{j}$ is already a systematic column in $\boldsymbol{G}_{i}$, so it remains in $\boldsymbol{G}_{i+1}$. Proceeding similarly as in Case 1, all service configurations in $\mathcal{S}_{i}$ remain feasible in $\mathcal{S}_{i+1}$.
Thus, $\mathcal{S}_{i}\subseteq\mathcal{S}_{i+1}$.
Strict Inclusion. To see that $\mathcal{S}_{i}$ is strictly contained in $\mathcal{S}_{i+1}$, consider the new systematic column $\boldsymbol{e}_{i+1}$, which is present in $\boldsymbol{G}_{i+1}$ but not in $\boldsymbol{G}_{i}$. From Proposition 3, the maximum single-basis service rate for $\boldsymbol{e}_{i+1}$ when it is systematic is
$\lambda_{i+1}^{\max}(\boldsymbol{G}_{i+1})\;=\;1+\frac{n-1}{k}\;>\;\frac{n}{k}\;=\;\lambda_{i+1}^{\max}(\boldsymbol{G}_{i}),\,\text{for all }k\geq 2.$
[…]
Figure 4: Rate polytopes $\mathcal{S}_{i}(12,3)$ for $i=0,1,2,3$. As $i$ increases, the SRR region strictly expands.
Figure 4 demonstrates this increasing inclusion for cases $n=12$ and $k=3$. We observe that increasing the number of systematic columns expands the service region in storage systems employing MDS codes with the same number of servers. The list of linear constraints that characterize these regions are derived in Section VI.
V Greedy Matching Allocation
This section introduces an efficient allocation scheme called Greedy matching, inspired by the classical Greedy algorithm. The key idea is to serve each data request using its associated systematic server as much as possible, since systematic servers store uncoded copies of data and can independently provide the requested data object without collaboration with other servers.
We then prove that for any system employing an MDS code generated by a matrix $\boldsymbol{G}$ and a service rate $\boldsymbol{\lambda}\in\mathcal{S}(\boldsymbol{G})$, there exists a matching $\boldsymbol{w}$ such that $\boldsymbol{\lambda}=\boldsymbol{w}\boldsymbol{S}$ and $\boldsymbol{w}$ can be realized through Greedy matching, as defined below. This allocation method will be critical in characterizing the service regions of MDS codes in the next section.
Definition 3 (Greedy Matching for MDS-coded Systems).
Consider any vector $\boldsymbol{\lambda}=(\lambda_{1},\ldots,\lambda_{k})\in\mathcal{S}_{i}$. The Greedy matching scheme allocates as much of each request $\lambda_{j}$ as possible to its systematic recovery set (i.e., systematic server), while any remaining portion is assigned to non-systematic recovery sets. Concretely, for each request $\lambda_{j}$, $j\in[i]$, define:
$\lambda_{j}^{\textsf{s}}\;=\;\min\{\lambda_{j},\,1\}\quad\text{and}\quad\lambda_{j}^{\textsf{c}}\;=\;\lambda_{j}-\lambda_{j}^{\textsf{s}},$
where $\lambda_{j}^{\textsf{s}}$ and $\lambda_{j}^{\textsf{c}}$ denote, respectively, the fraction of $\lambda_{j}$ served by the systematic (“s”) and non-systematic (“c”) servers. Consequently, all remaining requests
$\sum_{j=1}^{i}\lambda_{j}^{\textsf{c}}\;+\;\sum_{j=i+1}^{k}\lambda_{j}\;=\;\sum_{j=1}^{i}\bigl{(}\lambda_{j}-1\bigr{)}^{+}\;+\;\sum_{j=i+1}^{k}\lambda_{j}$
will be served by non-systematic nodes and any unused portion of the systematic nodes.
Theorem 2.
Any $\boldsymbol{\lambda}\in\mathcal{S}(\boldsymbol{G})$ can be served via Greedy matching.
Proof.
Let $\boldsymbol{G}=\boldsymbol{G}_{i}(n,k)$ be fixed, and denote $\Gamma_{G}$ as its associated recovery hypergraph. Because $\boldsymbol{\lambda}$ is servable by $\boldsymbol{G}$, by Proposition 1 there exists a matching $M_{1}$ in $\Gamma_{G}$ such that the total weight of the hyperedges labeled by $\boldsymbol{e}_{l}$ is $\lambda_{l}$ for each $l\in[k]$, i.e., $\boldsymbol{\lambda}(M_{1})=M_{1}\boldsymbol{S}$.
Suppose that $M_{1}$ is not greedy, i.e., there is some $l\in[i]$ for which the portion of $\lambda_{l}$ served by its systematic column is
$\lambda_{l}^{\mathsf{s}}(M_{1})\;<\;\min\{1,\;\lambda_{l}\}.$
We will construct another matching $M_{2}$ with
$\begin{cases}\lambda_{l}^{\mathsf{s}}(M_{2})\;>\;\lambda_{l}^{\mathsf{s}}(M_{1}),\\ \lambda_{j}^{\mathsf{s}}(M_{2})\;=\;\lambda_{j}^{\mathsf{s}}(M_{1}),&\forall\,j\neq l,\ j\in[i].\end{cases}$ (12)
We consider two cases:
Case 1: The systematic server $G^{l}=\boldsymbol{e}_{l}$ is not saturated.
Because $\lambda_{l}^{\mathsf{s}}(M_{1})<\min\{1,\lambda_{l}\}$, some of $\lambda_{l}$ is allocated to non-systematic edges instead. Since $G^{l}$ is not fully utilized, we can shift a small portion $\delta>0$ of $\lambda_{l}$ onto $G^{l}$. On the other hand, choose any hyperedge serving $\lambda_{l}^{\mathsf{c}}$ and reduce its weight by $\delta$. Because $G^{l}$ was not saturated, no node in the system becomes overused after the shift, preserving feasibility. Thus, we obtain a valid matching $M_{2}$ with
$\lambda_{l}^{\mathsf{s}}(M_{2})\;=\;\lambda_{l}^{\mathsf{s}}(M_{1})+\delta\;>\;\lambda_{l}^{\mathsf{s}}(M_{1}),$
and $\lambda_{j}^{\mathsf{s}}(M_{2})=\lambda_{j}^{\mathsf{s}}(M_{1})$ for all $j\neq l$.
Case 2: The systematic column $G^{l}$ is saturated.
Although $G^{l}$ is fully occupied, the condition $\lambda_{l}^{\mathsf{s}}(M_{1})<\min\{1,\lambda_{l}\}$ implies that some portion of $G^{l}$ is being used to recover a different vector $\boldsymbol{e}_{m}$ with $m\neq l$. Let $\epsilon$ be a non-systematic edge containing $G^{l}$ and labeled by $\boldsymbol{e}_{m}$, and let $w=\lambda_{m,\epsilon}$ be its weight.
Similarly, the condition $\lambda_{l}^{\mathsf{s}}(M_{1})<\min\{1,\lambda_{l}\}$ also implies the existence of a non-systematic edge $\epsilon_{1}$ (i.e., $\epsilon_{1}\not\ni G^{l}$) that contributes to the recovery of $\lambda_{l}$. Let its weight be $\lambda_{l,\epsilon_{1}}=w_{1}$.
Since $\epsilon$ and $\epsilon_{1}$ differ in at least one node—one contains $G^{l}$, while the other does not—it follows that $|\epsilon\cup\epsilon_{1}|\geq k+1$.
  1. 1.
    If $w_{1}\leq w$, then we reduce by $w_{1}$ the weight on both $\epsilon$ and $\epsilon_{1}$. We reassign that released $w_{1}$ portion of $G^{l}$ to $\lambda_{l}$. Simultaneously, we use a newly freed $k$-subset of nodes (excluding $G^{l}$, we can do that since $|\epsilon\cup\epsilon_{1}|\geq k+1$) to serve $\boldsymbol{e}_{m}$. This yields a valid matching $M_{2}$ satisfying (12).
  2. 2.
    If $w_{1}>w$, we release $w$ amount from $\epsilon$ and $\epsilon_{1}$ instead, then apply the same argument.
In either case, we can incrementally increase $\lambda_{l}^{\mathsf{s}}(M)$ while maintaining a valid matching. Repeating this argument shows that we can continue shifting capacity until
$\lambda_{l}^{\mathsf{s}}(M_{s})\;=\;\min\{1,\;\lambda_{l}\}$
for some matching $M_{s}$. Thus, all of $\lambda_{l}$ is served systematically (if $\lambda_{l}\leq 1$) or the systematic node is fully saturated by $\lambda_{l}$ (if $\lambda_{l}\geq 1$).
Hence, any achievable service vector $\boldsymbol{\lambda}$ under some allocation can be served by a Greedy matching that prioritizes each systematic column up to $\min\{\lambda_{l},1\}$. This proves the claimed result: sending each request to its systematic node until it is either served entirely or that node is saturated does not compromise feasibility. ∎ Mathematically, Greedy matching allows us to pre-assign values to certain edge weights in the vector $\boldsymbol{w}$ before checking the feasibility of the linear program in (4). This pre-assignment substantially reduces the degrees of freedom, significantly reducing the feasibility check’s complexity. Concretely, for a recovery hypergraph $\Gamma_{i}$ of $\boldsymbol{G}_{i}$ and a demand $\lambda_{j}$ such that $j\leq i$ (i.e., $\boldsymbol{G}_{i}$ contains a systematic node for object $j$):
  • If $\lambda_{j}\geq 1$, then a weight $w=1$ is assigned to the systematic edge labeled $\boldsymbol{e}_{j}$. Since the corresponding server is fully utilized, all other recovery sets that contain this node $\boldsymbol{e}_{j}$ are assigned weight 0 and thus can be removed.
  • If $\lambda_{j}<1$, then a weight $w=\lambda_{j}$ is assigned to the systematic edge labeled $\boldsymbol{e}_{j}$. Since the demand for object $j$ is fully served, all recovery sets labeled by $\boldsymbol{e}_{j}$ are assigned weight 0 and thus can be removed.
This process is demonstrated in Figure 5. Furthermore, in Theorem 7 in the final section, we show cases where Greedy matching is the only possible rate-splitting scheme to achieve certain request vectors in the SRR.
Remark 5.
A special case of this theorem, where $\boldsymbol{G}$ is systematic, was proven in [2], Lemmas 1 and 2. In this work, the authors established that it is optimal first to send requests to their corresponding systematic node. Based on this result, they devised the water-filling Algorithm for request rate splitting. The high-level idea behind this algorithm is that requests are first routed to the respective systematic (uncoded) server. Once the systematic servers are saturated, the requests are sent to the $k$ currently least-loaded servers, which can collaboratively form non-systematic recovery sets.
Figure 5: Recovery graph $\Gamma_{2}(4,2)$ (left) and Greedy matching on $\Gamma_{2}(4,2)$ when $\lambda_{1}\geq 1$ and $\lambda_{2}<1$ (right). After the Greedy matching, the systematic edge labeled $\boldsymbol{e}_{2}$ is assigned a weight of $\lambda_{2}<1$, while all other edges labeled $\boldsymbol{e}_{2}$ are assigned a weight of 0 and thus removed. The systematic edge labeled $\boldsymbol{e}_{1}$ is assigned a weight of 1. The degrees of freedom, which equal the number of edges without a weight, are reduced from 8 to 3.
VI Characterizing the SRR Polytopes
In this section, we develop a linear characterization of the SRRs for MDS-coded systems. We begin by recalling that $\mathcal{S}_{0}$, the SRR polytope of $\boldsymbol{G}_{0}(n,k)$, is described by the following $k+1$ linear constraints:
$\begin{cases}\sum_{j=1}^{k}\lambda_{j}\;\leq\;{n}/{k},&\quad(\text{1 constraint}),\\[4.0pt] \lambda_{j}\;\geq\;0,\quad\forall\,j\in[k],&\quad(k\text{ constraints}).\end{cases}$
The service polytope of $\boldsymbol{G}_{0}(6,3)$, which is a simplex, is shown in Fig. 6.
$\lambda_{1}$$\lambda_{2}$$\lambda_{3}$$222$
Figure 6: The service polytope of $\boldsymbol{G}_{0}(6,3)$.
VI-A Bounding the SRR via Fractional Matchings
We now generalize this description to the SRRs corresponding to other generator matrices $\boldsymbol{G}_{i}(n,k)$. We begin by recalling a useful lemma from [13]:
Lemma 1 ([13], Lemma IV.1).
For any fractional matching $\boldsymbol{w}\in\mathbb{R}^{|E|}$ on a hypergraph $(V,E)$, the following inequality holds:
$|V|\;\geq\;\sum_{\epsilon\in E}w_{\epsilon}\,|\epsilon|.$
This inequality is referred to as the capacity bound, where the left-hand side represents the total capacity of all nodes. A matching that satisfies this bound with equality is called perfect.
This result helps characterize the SRR polytopes of MDS codes described in Section IV. Specifically, by leveraging the Greedy matching approach (Section V), we show that the set of achievable demand vectors $\boldsymbol{\lambda}\in\mathcal{S}_{i}$ must lie within certain linear bounds derived from Lemma 1. We then prove that these bounds can be attained in many cases, thereby fully characterizing the corresponding SRR polytopes.
Greedy Allocation and Fractional-Matching Bound. Consider a coding scheme defined by $\boldsymbol{G}_{i}(n,k)$, and let $\boldsymbol{\lambda}=(\lambda_{1},\dots,\lambda_{k})\in\mathcal{S}_{i}$ be any achievable demand vector. Without loss of generality, assume $\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{k}$. Define $i_{A}$ as the unique number in $[i]$ so that
$\lambda_{j}\;\geq\;1\quad\text{for all }j\leq i_{A},\quad\text{and}\quad\lambda_{j}\;<\;1\quad\text{for }i_{A}<j\leq i.$
Under the Greedy matching scheme (see Section V), we allocate:
  1. 1.
    $\lambda_{j}^{\mathsf{s}}=1$ for each $j\leq i_{A}$,
  2. 2.
    $\lambda_{j}^{\mathsf{s}}=\lambda_{j}$ for each $i_{A}<j\leq i$,
  3. 3.
    $\lambda_{j}^{\mathsf{s}}=0$ for each $j>i$.
Hence, in the greedy matching step, each systematic server associated with a basis vector $\boldsymbol{e}_{j}$, $j\leq i_{A}$ is fully saturated, while servers associated with basis vectors $\boldsymbol{e}_{j}$, $i_{A}<j\leq i$ are utilized up to $\lambda_{j}$, where $\lambda_{j}<1$. All remaining demands
$\sum_{j=1}^{i_{A}}(\lambda_{j}-1)\;+\;\sum_{j=i+1}^{k}\lambda_{j}$
must be served by non-systematic edges of size $k$. By Lemma 1, these non-systematic edges impose the following constraint:
$\left(i_{A}+\sum_{j=i_{A}+1}^{i}\lambda_{j}\right)+k\cdot\left(\sum_{j=1}^{i_{A}}(\lambda_{j}-1)\;+\;\sum_{j=i+1}^{k}\lambda_{j}\right)\;\;\leq\;\;n,$
which can be written as:
$k\sum_{j=1}^{i_{A}}\lambda_{j}\;+\;\sum_{j=i_{A}+1}^{i}\lambda_{j}\;+\;k\sum_{j=i+1}^{k}\lambda_{j}\;\leq\;n+k-i_{A}.$ (13)
We now see that this bound can be tight, as the next result shows.
Theorem 3 (Subgraph Slicing).
If $n-i\geq k$ then the bound in (13) is achievable.
Proof.
We proceed through the following steps.
Post-Greedy Set-Up. Under Greedy Matching, each systematic server associated with a basis vector $\boldsymbol{e}_{j}$, $j\leq i_{A}$ is fully saturated. Those associated with $\boldsymbol{e}_{j}$, $i_{A}<j\leq i$ were partially used, leaving capacity $1-\lambda_{j}$. Finally, any node beyond node $i$ remains fully available (capacity $1$). Formally, the capacity of node $G^{j}$ after Greedy is
$\text{Cap}(G^{j})\;=\;\begin{cases}0,&j\leq i_{A},\\ 1-\lambda_{j},&i_{A}<j\leq i,\\ 1,&j>i.\end{cases}$
Because $n-i\geq k$, at least $k$ columns (nodes) remain entirely unused, which will be crucial to constructing further $k$-sized matchings.
Residual Hypergraph H. Let $H$ denote the residual hypergraph after the Greedy step, where each node has the above capacity. We want to show that the leftover demands
$\sum_{j=1}^{i_{A}}(\lambda_{j}-1)\;+\;\sum_{j=i+1}^{k}\lambda_{j}$
are exactly matched by size-$k$ edges in $H$. However, node capacities in $H$ are non-uniform: some have capacity $1-\lambda_{j}$, others have capacity $1$. To handle this, we “slice” $H$ into smaller $k$-uniform hypergraphs whose nodes each have a uniformly assigned capacity.
Slicing Construction. We form $(i-i_{A}+1)$ sub-hypergraphs $H_{1},\dots,H_{i-i_{A}+1}$, each resembling a $k$-uniform hypergraph on some subset of columns:
  • $H_{1}$ uses $(n-i_{A})$ columns each at capacity $\alpha_{1}=1-\lambda_{i_{A}+1}$ (or $\alpha_{1}=1$ if $i_{A}=i$).
  • $H_{2}$ has $(n-i_{A}-1)$ columns each at capacity $\alpha_{2}=\lambda_{i_{A}+1}-\lambda_{i_{A}+2}$, and so on.
  • In general, each $H_{\ell}$ has uniform capacity $\alpha_{\ell}=\lambda_{i_{A}+\ell-1}-\lambda_{i_{A}+\ell}$ among a progressively smaller set of columns.
  • $H_{\,i-i_{A}+1}$ (the final slice) has $\alpha_{\,i-i_{A}+1}=\lambda_{i}$ among $(n-i)$ columns.
(If some $\lambda_{j}$ are zero for $j>i$ or $j\leq i_{A}$, we adjust accordingly, but the concept remains the same.)
Capacity Summation. Since $n-i_{A}\geq n-i\geq k$, each $H_{\ell}$ is essentially a $k$-uniform hypergraph on $m_{\ell}$ vertices (columns), where $m_{\ell}$ denotes the number of vertices in $H_{\ell}$ and satisfies $m_{\ell}\geq k$. By the known $\mathcal{S}_{0}$ bound $\sum_{j=1}^{k}\lambda_{j}\leq\tfrac{m_{\ell}}{k}$, we can serve exactly $\alpha_{\ell}\,\tfrac{m_{\ell}}{k}$ units of demand from the slice $H_{\ell}$. Summing over slices, the total capacity becomes
$(1-\lambda_{i_{A}+1})\dfrac{n-i_{A}}{k}+(\lambda_{i_{A}+1}-\lambda_{i_{A}+2})\dfrac{n-i_{A}-1}{k}+\cdots+(\lambda_{i-1}-\lambda_{i})\dfrac{n-i+1}{k}+\lambda_{i}\dfrac{n-i}{k}$
$=\dfrac{n-i_{A}}{k}-\dfrac{1}{k}\sum\limits_{j=i_{A}+1}^{i}\lambda_{j}.$
Matching the Leftover Demands Exactly. The leftover demand after Greedy is
$\sum_{j=1}^{i_{A}}(\lambda_{j}-1)\;+\;\sum_{j=i+1}^{k}\lambda_{j}.$
Because the slices $H_{\ell}$ are $k$-uniform sub-hypergraphs with uniform capacity $\alpha_{\ell}$, each can form a fractional (or integral) matching that fully utilizes its $\alpha_{\ell}\,\tfrac{m_{\ell}}{k}$ capacity. By choosing appropriate edges (of size $k$) in each sub-hypergraph, we allocate these capacities to meet the leftover demands perfectly. Hence, equality
$\sum_{j=1}^{i_{A}}(\lambda_{j}-1)\;+\;\sum_{j=i+1}^{k}\lambda_{j}=\dfrac{n-i_{A}}{k}-\dfrac{1}{k}\sum\limits_{j=i_{A}+1}^{i}\lambda_{j}.$
can be attained in the residual hypergraph $H$. In other words, (13) can be achieved with equality.
Conclusion. Thus, by constructing the sub-hypergraphs $H_{1},\dots,H_{\,i-i_{A}+1}$ and allocating the leftover demands slice by slice, we show that the bound in Eq. (13) is exactly achievable when $n-i\geq k$. This completes the proof. ∎
Figure 7: Illustration of a storage system $\boldsymbol{G}_{3}(6,3)$. In this example, $\lambda_{1}>1$, $\lambda_{2}=0.8$, and $\lambda_{3}=0.3$, so $i_{A}=1$. After Greedy matching, the hypergraph $H$ is “sliced” into three smaller $3$-uniform hypergraphs: the first (green) spans the last five vertices with $\alpha_{1}=1-\lambda_{2}=0.2$, the second (purple) spans the last four vertices with $\alpha_{2}=\lambda_{2}-\lambda_{3}=0.5$, and the third (red) spans the last three vertices with $\alpha_{3}=\lambda_{3}=0.3$. The maximum additional request for $\lambda_{1}$ that the system can serve after the Greedy matching is therefore constrained by $\lambda_{1}-1\leq\alpha_{1}\cdot 5/3+\alpha_{2}\cdot 4/3+\alpha_{3}\cdot 3/3=1.3$.
Figure 7 illustrates this subgraph slicing process based on the values of $\lambda_{1},\lambda_{2},$ and $\lambda_{3}$.
Remark 6.
Equation (13) can be equivalently rewritten as
$\sum_{j=1}^{i_{A}}\lambda_{j}\;+\;\sum_{j=i+1}^{k}\lambda_{j}\;-\;i_{A}\;\;=\;\;\frac{n-i_{A}-\sum_{j=i_{A}+1}^{i}\lambda_{j}}{k}.$ (14)
When $i_{A}=k$ (meaning $\boldsymbol{G}=\boldsymbol{G}_{k}$ is a systematic matrix), this recovers Theorem 1 of [2] as a special case. In other words, (14) generalizes that result to systems where some (but not all) columns are systematic, showing that similarly tight bounds hold for both systematic and partially systematic codes.
We move on to characterize the SRRs in some specific cases.
VI-B $\mathcal{S}_{i}(n,k)$ when $n\geq k+i$
In this scenario, $n-i\geq k$. By Theorem 3, the bound therein is achieved, implying
$k\Bigl{(}\sum_{j=1}^{i_{A}}\lambda_{j}+\sum_{j=i+1}^{k}\lambda_{j}\Bigr{)}\;+\;\sum_{j=i_{A}+1}^{i}\lambda_{j}\;=\;n+i_{A}(k-1).$ (15)
Let us define a partition of the index set $\{1,2,\ldots,i\}$ into three subsets $\mathcal{\mathcal{missing}}{A},\mathcal{\mathcal{missing}}{B},\mathcal{\mathcal{missing}}{C}$ as follows:
$\mathcal{\mathcal{missing}}{A}\;=\;\{\,j\in[i]:\lambda_{j}\geq 1\},\qquad\mathcal{\mathcal{missing}}{B}\;=\;[i]\setminus\mathcal{\mathcal{missing}}{A},\qquad\mathcal{\mathcal{missing}}{C}\;=\;\{\,i+1,i+2,\dots,k\}.$
Then (15) can be rewritten as
$k\Bigl{(}\sum_{j\in\mathcal{\mathcal{missing}}{A}}\lambda_{j}+\sum_{j\in\mathcal{\mathcal{missing}}{C}}\lambda_{j}\Bigr{)}\;+\;\sum_{j\in\mathcal{\mathcal{missing}}{B}}\lambda_{j}\;=\;n\;+\;|\mathcal{\mathcal{missing}}{A}|\,(k-1).$
Because any choice of $\mathcal{\mathcal{missing}}{A}\subseteq[i]$ uniquely determines $\mathcal{\mathcal{missing}}{B}=[i]\setminus\mathcal{\mathcal{missing}}{A}$, and there are $2^{i}$ ways to pick $\mathcal{\mathcal{missing}}{A}\subseteq[i]$, we obtain a family of linear constraints describing $\mathcal{S}_{i}$. Concretely, for $\boldsymbol{G}_{i}(n,k)$ with $n\geq k+i$, the SRR $\mathcal{S}_{i}$ satisfies:
Theorem 4.
If $n\geq k+i$, the service region $\mathcal{S}_{i}$ is given by the set of all nonnegative $\boldsymbol{\lambda}$ satisfying:
$\begin{cases}\lambda_{l}\geq 0,&\forall\,l\in[k],\\[4.0pt] k\,\Bigl{(}\sum_{j=1}^{k}\lambda_{j}\Bigr{)}\;\leq\;n+i\,(k-1),&\\[4.0pt] \lambda_{l}\;+\;k\Bigl{(}\sum_{j\in[k]\setminus\{l\}}\lambda_{j}\Bigr{)}\;\leq\;n\;+\;(i-1)\,(k-1),&\forall\,l\in[i],\\[4.0pt] \lambda_{l}\;+\;\lambda_{h}\;+\;k\Bigl{(}\sum_{j\in[k]\setminus\{l,h\}}\lambda_{j}\Bigr{)}\;\leq\;n\;+\;(i-2)\,(k-1),&\forall\,l,h\in[i],\;l\neq h,\\[2.0pt] \quad\vdots&\\[2.0pt] \sum_{l\in[i]}\lambda_{l}\;+\;k\Bigl{(}\sum_{j\in[k]\setminus[i]}\lambda_{j}\Bigr{)}\;\leq\;n.&\end{cases}$ (16)
There are $k+2^{i}$ constraints in total: $k$ nonnegativity constraints and $2^{i}$ constraints arising from all subsets $\mathcal{\mathcal{missing}}{A}\subseteq[i]$. In the special case $i=k$, the second and last constraints in (16) become linearly dependent. Thus, only the tighter one—the last constraint—remains effective, resulting in $12^{k}+k-$ 1 distinct constraints.
Remark 7.
If $n\geq 2k$, the condition $n\geq k+i$ is always satisfied since $i\leq k$. Therefore, for any MDS-coded system with at least twice as many servers as data objects, the SRR is fully determined regardless of the number of systematic servers.
VI-C $\mathcal{S}_{k}(k+1,k)$ of systematic coding
Now consider the matrix $\boldsymbol{G}=\boldsymbol{G}_{k}(k+1,k)$, which corresponds to a systematic code with $k$ systematic columns and 1 parity column. Hence, $\mathcal{\mathcal{missing}}{C}=\varnothing$ and the partition reduces to $\mathcal{\mathcal{missing}}{A},\mathcal{\mathcal{missing}}{B}\subseteq[k]$ with $\mathcal{\mathcal{missing}}{A}\cup\mathcal{\mathcal{missing}}{B}=[k]$. Then (13) simplifies to
$k\,\Bigl{(}\sum_{j\in\mathcal{\mathcal{missing}}{A}}\lambda_{j}\;-\;|\mathcal{\mathcal{missing}}{A}|\Bigr{)}\;+\;\sum_{j\in\mathcal{\mathcal{missing}}{B}}\lambda_{j}\;\;\leq\;\;(k+1)\;-\;|\mathcal{\mathcal{missing}}{A}|.$ (17)
We show that in this case, each pair of distinct requests $\lambda_{i},\lambda_{j}$ must satisfy $\lambda_{i}+\lambda_{j}\leq 2$. Formally:
Lemma 2.
For any data request $\boldsymbol{\lambda}\in\mathcal{S}_{k}(k+1,k)$, one has
[…]
Proof.
In the recovery graph $\Gamma_{G}$, consider the subgraph $\Gamma^{\prime}$ obtained by removing all edges except those labeled by $\boldsymbol{e}_{i}$ and $\boldsymbol{e}_{j}$. Assign weight 1 to the two nodes corresponding to the systematic columns for $i$ and $j$, and weight 0 to all other nodes. Since the remaining $(k-1)$ nodes cannot independently form a recovery set for either $\boldsymbol{e}_{i}$ or $\boldsymbol{e}_{j}$ without involving their respective systematic columns, this weight assignment forms a valid vertex cover of $\Gamma^{\prime}$ of size 2. Consequently, Proposition 2 implies $\lambda_{i}+\lambda_{j}\leq 2$. ∎
We then argue that this pairwise constraint, along with (17) and nonnegativity, fully characterizes the SRR polytope:
Theorem 5.
The service region $\mathcal{S}_{k}(k+1,k)$ is given by
[…]
$\lambda_{i}+\lambda_{j}$  $\;\leq\;2,\quad\forall\,i\neq j,\;i,j\in[k],$ (vertex-cover constraints) (19)
$k\,\Bigl{(}\sum_{j\in\mathcal{\mathcal{missing}}{A}}\lambda_{j}\;-\;|\mathcal{\mathcal{missing}}{A}|\Bigr{)}+\sum_{j\in\mathcal{\mathcal{missing}}{B}}\lambda_{j}$ $\;\leq\;k+1-|\mathcal{\mathcal{missing}}{A}|,$ (node-capacity constraint) (20)
where $\{\mathcal{\mathcal{missing}}{A},\mathcal{\mathcal{missing}}{B}\}$ is a partition of $[k]$ and $\mathcal{\mathcal{missing}}{A}$ is the set of indices $j$ with $\lambda_{j}>1$.
Proof.
Because $\boldsymbol{G}$ includes systematic columns for all $k$ objects, the roles of $\lambda_{1},\dots,\lambda_{k}$ are symmetric. Without loss of generality, let $\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{k}$. Consider the constraint $\lambda_{1}+\lambda_{2}\leq 2$, which implies that only $\lambda_{1}$ can exceed 1.
Case 1: $\lambda_{1}\leq 1$.
In this case, $\lambda_{j}\leq 1$ for all $j$. This request is clearly servable via Greedy matching, i.e., each request is served by its corresponding systematic server.
Case 2: $\lambda_{1}>1$.
Let $\lambda_{1}=1+\delta$ with $\delta>0$. From $\lambda_{1}+\lambda_{2}\leq 2$, it follows that $\lambda_{2}\leq 1-\delta$. Thus, we have $\lambda_{1}>1>\lambda_{2}\geq\dots\geq\lambda_{k}$.
After Greedy matching, requests $\lambda_{2},\dots,\lambda_{k}$ are fully served by their corresponding systematic columns. The remaining portion, $\lambda_{1}-1=\delta$, can be served using a non-systematic recovery set consisting of:
  • the free capacity in $(k-1)$ systematic nodes $\boldsymbol{e}_{2},\dots,\boldsymbol{e}_{k}$ (each with at least $\delta$ available), and
  • the single unused parity column $p_{n}$ in $\boldsymbol{G}_{k}$.
Thus, $\lambda_{1}$ is fully servable.
These arguments show that (18), (19), and (20) indeed describe $\mathcal{S}_{k}(k+1,k)$ exactly. ∎
VI-D $\mathcal{S}_{i}(n,k)$ when $n=k+i-1$
In this case $\boldsymbol{G}=\boldsymbol{G}_{i}(k+i-1,k),\mathcal{C}=[k]\setminus[i]$ and $\mathcal{A}\cup\mathcal{B}=[i]$, therefore (14) becomes:
$k\Bigl{(}\sum\limits_{j\in\mathcal{A}\cup\mathcal{C}}\lambda_{j}-|\mathcal{A}|\Bigr{)}+\sum\limits_{j\in\mathcal{B}}\lambda_{j}$ $\leq k+i-1-|\mathcal{A}|$ (21)
This corner case is substantially more complicated than the previous cases when $n\geq k+i$.
We present and prove the following constraint:
Lemma 3.
For any service vector $\boldsymbol{\lambda}\in\mathcal{S}_{i}(k+i-1,k)$:
$\sum\limits_{j\in\mathcal{A}\cup\mathcal{C}}\lambda_{j}+\sum\limits_{j\in\mathcal{B}}\lambda_{j}=\sum\limits_{j=1}^{k}\lambda_{j}\leq i$ (22)
Proof.
Put a weight one on $i$ systematic vertices of the recovery graph $\Gamma_{G}$; the remaining $(k-1)$ nodes cannot independently form a non-systematic recovery set for any object without involving their respective systematic columns. Thus, we have a valid vertex cover of size $i$, by Proposition 2, $i\geq\sum\limits_{j=1}^{k}\lambda_{j}$. ∎
We will prove that this constraint and constraint (21), along with non-negativity constraints characterize the SRR polytope.
Theorem 6.
$\mathcal{S}_{i}(k+i-1,k)$ is given by:
$\lambda_{j}$ $\geq 0,\quad\forall j\in[k]$ (Non-negativity constraints) (23)
$\sum\limits_{j=1}^{k}\lambda_{j}$ $\leq i$ (Vertex cover constraint) (24)
$k\sum\limits_{j\in\mathcal{A}}(\lambda_{j}-1)+k\sum\limits_{j\in\mathcal{C}}\lambda_{j}+\sum\limits_{j\in\mathcal{B}}\lambda_{j}$ $\leq k+i-1-|\mathcal{A}|$ (Node capacity) (25)
Proof.
Let $a=|\mathcal{A}|$. For any non-negative request vector $\boldsymbol{\lambda}=(\lambda_{1},\lambda_{2},\ldots,\lambda_{k})$ that satisfies (24) and (25), we prove that it is servable by the system by considering one of the following 3 cases:
  1. 1.
    The Vertex cover constraint satisfied with equality:
    $\begin{cases}&k\sum\limits_{j\in\mathcal{A}}(\lambda_{j}-1)+k\sum\limits_{j\in\mathcal{C}}\lambda_{j}+\sum\limits_{j\in\mathcal{B}}\lambda_{j}\leq k+i-1-a\\ &\sum\limits_{j=1}^{k}\lambda_{j}=i\end{cases}$ (26)
    We first serve this request $\boldsymbol{\lambda}$ using Greedy matching. After Greedy matching allocation, the remaining request to be served is $\sum\limits_{i\in\mathcal{A}}(\lambda_{i}-1)+\sum\limits_{j\in\mathcal{C}}\lambda_{j}$. Note that because $n=k+i-1$, $k-1$ columns (nodes) remain entirely unused.
    We now prove that $\sum\limits_{j\in\mathcal{A}}(\lambda_{j}-1)+\sum\limits_{j\in\mathcal{C}}\lambda_{j}\leq 1$, i.e., $\sum\limits_{j\in\mathcal{A}\cup\mathcal{C}}\lambda_{j}\leq a+1$. Indeed, assume otherwise that $\sum\limits_{j\in\mathcal{A}\cup\mathcal{C}}\lambda_{j}=a+1+\delta$ for some $\delta>0$. From (26), we have
    $k+i-1-a$ $\geq k\sum\limits_{j\in\mathcal{A}}(\lambda_{j}-1)+k\sum\limits_{j\in\mathcal{C}}\lambda_{j}+\sum\limits_{j\in\mathcal{B}}\lambda_{j}$
    $=k\Bigl{(}\sum\limits_{j\in\mathcal{A}\cup\mathcal{C}}\lambda_{j}-a\Bigr{)}+\sum\limits_{j\in\mathcal{B}}\lambda_{j}$
    $=k\delta+k+\Bigl{(}i-\sum\limits_{j\in\mathcal{A}\cup\mathcal{C}}\lambda_{j}\Bigr{)}$
    $=k\delta+k+i-(a+1+\delta)$ (27)
    where the second equality comes from $i=\sum_{j=1}^{k}\lambda_{j}=\sum\limits_{j\in\mathcal{A}\cup\mathcal{C}}\lambda_{j}+\sum\limits_{j\in B}\lambda_{j}$. The last equation leads us to $0\geq k\delta-\delta$ or $1\geq k$, which could not happen when $k\geq 2$. Thus by contradiction, we have: $\sum\limits_{j\in\mathcal{A}\cup\mathcal{C}}\lambda_{j}\leq a+1$.
    Therefore, note that $\mathcal{A}\cup\mathcal{B}=[i]$, we have
    $1\geq\sum\limits_{j\in{\mathcal{A}\cup\mathcal{C}}}\lambda_{j}-a=i-\sum\limits_{j\in\mathcal{B}}\lambda_{j}-a=|\mathcal{B}|-\sum\limits_{j\in\mathcal{B}}\lambda_{j}=\sum\limits_{j\in\mathcal{B}}(1-\lambda_{j}).$
    which means that the sum of the free capacity of nodes in $\mathcal{B}$ after the greedy matching step is at most 1. Under this condition, we prove that the request rate can be served using the ensuing lemma.
    Lemma 4.
    (Successive Tiling) Under the coding scheme $\boldsymbol{G}_{i}(k+i-1,k)$, any request vector $\boldsymbol{\lambda}=(\lambda_{1},\lambda_{2},\dots,\lambda_{k})$ that satisfies the following constraints is servable:
    $\begin{cases}k\sum\limits_{j\in\mathcal{A}}(\lambda_{j}-1)+k\sum\limits_{j\in\mathcal{C}}\lambda_{j}+\sum\limits_{j\in\mathcal{B}}\lambda_{j}\leq k+i-1-a,\\ \sum\limits_{j\in\mathcal{A}\cup\mathcal{C}}\lambda_{j}+\sum\limits_{j\in\mathcal{B}}\lambda_{j}=\sum\limits_{j=1}^{k}\lambda_{j}=i,\\ \sum\limits_{j\in\mathcal{A}}(\lambda_{j}-1)+\sum\limits_{j\in\mathcal{C}}\lambda_{j}\leq 1.\end{cases}$ (28)
    Proof.
    We again serve this request using Greedy matching:
    $\begin{cases}\lambda_{j}^{\mathsf{s}}=1,\quad\forall j\in\mathcal{A},\\ \lambda_{j}^{\mathsf{s}}=\lambda_{j},\quad\forall j\in\mathcal{B},\\ \lambda_{j}^{\mathsf{s}}=0,\quad\forall j\in\mathcal{C}.\end{cases}$ (29)
    The remaining request to be served is:
    $\sum\limits_{j\in\mathcal{A}}(\lambda_{j}-1)+\sum\limits_{j\in\mathcal{C}}\lambda_{j}.$
    After the Greedy matching step, $k-1$ non-systematic nodes remain completely unused. For each systematic node $j\in\mathcal{B}$, we allocate a $(1-\lambda_{j})$ portion of all the $k-1$ non-systematic nodes. Together with the remaining $(1-\lambda_{j})$ portion of systematic node $j$, this forms a recovery set of size $k$ (a non-systematic recovery set), which can be used to serve any object with demand $1-\lambda_{j}$.
    From our assumption in (28), we have
    $\sum\limits_{j\in\mathcal{B}}(1-\lambda_{j})=\sum\limits_{j\in\mathcal{A}}(\lambda_{j}-1)+\sum\limits_{j\in\mathcal{C}}\lambda_{j}\leq 1.$
    Thus, we can perform this allocation for all systematic nodes $j\in\mathcal{B}$ without exhausting the $k-1$ non-systematic nodes. In the end, we have $|\mathcal{B}|$ non-systematic recovery sets of size $k$ with capacity $1-\lambda_{j}$ for $j\in\mathcal{B}$, which can be used to recover any data object.
    Since
    $\sum\limits_{j\in\mathcal{B}}(1-\lambda_{j})=\sum\limits_{j\in\mathcal{A}}(\lambda_{j}-1)+\sum\limits_{j\in\mathcal{C}}\lambda_{j},$
    all the remaining demand for objects in $\mathcal{A}$ and $\mathcal{C}$ can be served. This proves that $\boldsymbol{\lambda}$ is servable. ∎
[…]
    $\begin{cases}\mathcal{A}=\{1\},\,\mathcal{B}=\{2,3\},\,\mathcal{C}=[3]\setminus\{\mathcal{A}\cup\mathcal{C}\}=\varnothing\\ \sum\limits_{j\in\mathcal{A}\cup\mathcal{C}}\lambda_{j}+\sum\limits_{j\in\mathcal{B}}\lambda_{j}=3=i\end{cases}$
    Moreover, $\sum\limits_{j\in\mathcal{B}}(1-\lambda_{j})=1-0.6+1-0.5=0.9<1$.
    $\lambda_{1}$
    $\lambda_{2}$
    $\lambda_{3}$
    $\lambda_{1}$
    $\lambda_{2}$
    $\lambda_{3}$
[…]
  2. 2.
    The Node capacity constraint is tightly satisfied:
    $\begin{cases}&k\sum\limits_{j\in\mathcal{A}}(\lambda_{j}-1)+k\sum\limits_{j\in\mathcal{C}}\lambda_{j}+\sum\limits_{j\in\mathcal{B}}\lambda_{j}=k+i-1-a\\ &\sum\limits_{j=1}^{k}\lambda_{j}<i\end{cases}$ (30)
    Observe first that because $\sum\limits_{j=1}^{k}\lambda_{j}<i$, then there must exists $l\in[i]$ such that $\lambda_{l}<1$ (for otherwise $\sum\limits_{j=1}^{k}\lambda_{j}\geq\sum\limits_{j=1}^{i}\lambda_{j}\geq i$, contradiction!). Therefore, $a=|\mathcal{A}|\leq i-1$.
    We now find a lower bound for $\sum\limits_{j=1}^{k}\lambda_{j}$. Let $T=\sum\limits_{j\in\mathcal{B}}\lambda_{j}$, from (30) we have:
    $k\sum\limits_{j\in\mathcal{A}}(\lambda_{j}-1)+k\sum\limits_{j\in\mathcal{C}}\lambda_{j}$ $=k+i-1-a-T$ (31)
    $\Leftrightarrow\quad\sum\limits_{j=1}^{k}\lambda_{j}=\sum\limits_{j\in\mathcal{A}\cup\mathcal{C}}\lambda_{j}+T$ $=\dfrac{k+i-1-a-T}{k}+a+T$ (32)
    $=\dfrac{k+i-1-a}{k}+a+\Bigl{(}T-\dfrac{T}{k}\Bigr{)}$ (33)
    $\geq\dfrac{k+i-1-a}{k}+a$ (34)
    Therefore we see that $\Bigl{(}\dfrac{k+i-1-a}{k}+a\Bigr{)}$ is the lower bound for $\sum\limits_{j=1}^{k}\lambda_{j}$, and $\sum\limits_{j=1}^{k}\lambda_{j}=\dfrac{k+i-1-a}{k}+a$ if and only if $T=\sum\limits_{j\in\mathcal{B}}\lambda_{j}=0$ or
    $\begin{cases}{}&k\sum\limits_{j\in\mathcal{A}}(\lambda_{j}-1)+k\sum\limits_{j\in\mathcal{C}}\lambda_{j}=k+i-1-a\\ &\sum\limits_{j\in\mathcal{B}}\lambda_{j}=0\end{cases}$
    The second constraint $\sum\limits_{j\in\mathcal{B}}\lambda_{j}=0$ means that $\lambda_{j}=0$ for all $j\in\mathcal{B}$. Therefore, all the nodes in $\mathcal{B}$ are untouched, and the Node capacity constraint $k\sum\limits_{j\in\mathcal{A}}(\lambda_{j}-1)+k\sum\limits_{j\in\mathcal{C}}\lambda_{j}=k+i-1-a$ can be tightly satisfied just like the case $\boldsymbol{G}_{i}(n,k)$ when $n\geq k+i$ (recall that we have $|\mathcal{A}|=a\leq i-1$, so after the Greedy matching allocation, we will be left with at least $(k+i-1)-(i-1)=k$ entirely unused nodes).
    We have just proved that for any request vector $\boldsymbol{\lambda}_{A}$ such that:
    $\begin{cases}{}&k\sum\limits_{j\in\mathcal{A}}(\lambda_{j}-1)+k\sum\limits_{j\in\mathcal{C}}\lambda_{j}+\sum\limits_{j\in\mathcal{B}}\lambda_{j}=k+i-1-a\\ &\sum\limits_{j=1}^{k}\lambda_{j}=\dfrac{k+i-1-a}{k}+a\end{cases}$
    is servable. On the other hand, part 1 proved that any request vector $\boldsymbol{\lambda}_{B}$ such that
    $\begin{cases}{}&k\sum\limits_{j\in\mathcal{A}}(\lambda_{j}-1)+k\sum\limits_{j\in\mathcal{C}}\lambda_{j}+\sum\limits_{j\in\mathcal{B}}\lambda_{j}=k+i-1-a\\ &\sum\limits_{j=1}^{k}\lambda_{j}=i\end{cases}$
    is also servable. For any request vector $\boldsymbol{\lambda}$ that tightly satisfies the Node capacity constraint (i.e., the first part of (30) holds), its sum rate $\sum\limits_{j=1}^{k}\lambda_{j}$ must lie within the range
    $\Biggl{[}\frac{k+i-1-a}{k}+a,i\Biggr{]}.$
    Here, we have established that $\frac{k+i-1-a}{k}+a$ serves as the lower bound for $\sum\limits_{j=1}^{k}\lambda_{j}$, and Lemma 22 has proven that its upper bound is $i$. Consequently, $\boldsymbol{\lambda}$ must be servable, as it can be expressed as a linear combination of the two servable extreme points $\boldsymbol{\lambda}_{A}$ and $\boldsymbol{\lambda}_{B}$.
  3. 3.
    None of the two constraints is tightly satisfied:
    $\begin{cases}&k\sum\limits_{j\in\mathcal{A}}(\lambda_{j}-1)+k\sum\limits_{j\in\mathcal{C}}\lambda_{j}+\sum\limits_{j\in\mathcal{B}}\lambda_{j}<n-a=k+i-1-a\\ &\sum\limits_{j=1}^{k}\lambda_{j}<i\end{cases}$ (35)
    In this case, we can always increase 1 request component $\lambda_{j}$ until either the first or second part of (35) becomes equality (whichever comes first), and we’re back to the previous cases (1) and (2). Thus, any vector that falls into this case can be served.
Fig. 9 plots the SRRs of $\boldsymbol{G}_{3}(4,3)$ and $\boldsymbol{G}_{3}(5,3)$. The SRR of $\boldsymbol{G}_{3}(5,3)$ is given by:
$\begin{cases}\lambda_{j}\geq 0,\quad j\in[3],\\ \sum\limits_{j=1}^{3}\lambda_{j}\leq 3,&\\ 3\lambda_{1}+\lambda_{2}+\lambda_{3}\leq 7,\quad 3\lambda_{2}+\lambda_{1}+\lambda_{3}\leq 7,\quad 3\lambda_{3}+\lambda_{2}+\lambda_{1}\leq 7,&\\ 3(\lambda_{1}+\lambda_{2})+\lambda_{3}\leq 9,\quad 3(\lambda_{1}+\lambda_{3})+\lambda_{2}\leq 9,\quad 3(\lambda_{2}+\lambda_{3})+\lambda_{1}\leq 9.&\end{cases}$
Note that the constraint $\sum\limits_{j=1}^{3}\lambda_{j}\leq 3$ is tighter than the last three constraints, rendering them inactive and redundant.
$\lambda_{c}$$\lambda_{b}$$\lambda_{a}$$\lambda_{c}$$\lambda_{b}$$\lambda_{a}$
Figure 9: Service polytope of systematic MDS code $\boldsymbol{G}_{3}(4,3)$ (left) and $\boldsymbol{G}_{3}(5,3)$.
VI-E Coding schemes with $k\leq n<k+i-1$
In this case, the general SRR is unknown. However, we prove that although Lemma 22 still holds, it can not be satisfied by too many vectors $\boldsymbol{\lambda}$. We see from Fig. 9 that in $\mathcal{S}_{3}(5,3)$, there is a plane of request vectors $\boldsymbol{\lambda}$ such that $\sum_{j=1}^{k}\lambda_{j}=3$, while in $\mathcal{S}_{3}(4,3)$ there is only one such vector, $\boldsymbol{\lambda}=(1,1,1)$, that satisfies this constraint. In general, the SRR of $\boldsymbol{G}_{i}(k+i-1,k)$ contains a plane of service vectors $\boldsymbol{\lambda}$ such that $\sum\limits_{j=1}^{k}\lambda_{j}=i$.
We will prove that when $n<k+i-1$, there is only one vector $\boldsymbol{\lambda}$ that satisfies the constraint $\sum\limits_{j=1}^{k}\lambda_{j}=i$, namely $\boldsymbol{\lambda}=(1,1,\ldots,1,0,0,\ldots,0)$ (with the first $i$ elements equal to 1 and the remaining elements equal to 0). The key idea behind the proof is that when $n$ is too small relative to $k+i-1$, the number of non-systematic nodes is insufficient to support any other allocation satisfying the given constraint. This result shows that in this case, apart from Greedy matching, no other rate-splitting scheme can serve this request vector.
Theorem 7.
For $k\leq n\leq k+i-2$, the service polytope $\mathcal{S}_{i}(n,k)$ contains only the vector
$\begin{array}[]{c}\boldsymbol{\lambda}=(1,\ldots,1,0,\ldots,0)\vspace{-11mm}\\ \underbrace{\hskip 34.14322pt}_{\text{$i$ elements}}\end{array}$
as the unique vector satisfying $\sum\limits_{j=1}^{k}\lambda_{j}=i$. Moreover, any scheme other than Greedy matching cannot serve this request vector.
Proof.
Since $\Gamma_{G}$ has one systematic node for each of the first $i$ basis vectors, the vector $\boldsymbol{\lambda}$ is servable, meaning it lies within the service polytope. Assume, toward a contradiction, that there exists another vector $\boldsymbol{\mu}\neq\boldsymbol{\lambda}$ in the polytope such that $\sum\limits_{j=1}^{k}\mu_{j}=i$. Theorem 2 guarantees that this vector is servable by Greedy matching. Applying Greedy matching to $\boldsymbol{\mu}$, we obtain:
$\begin{cases}(\mu_{j})^{\mathsf{s}}=1,\quad\forall j\in\mathcal{A},\\ (\mu_{j})^{\mathsf{s}}=\mu_{j},\quad\forall j\in\mathcal{B},\\ (\mu_{j})^{\mathsf{s}}=0,\quad\forall j\in\mathcal{C}.\end{cases}$
After this step, the remaining request to be served is
$\sum\limits_{j\in\mathcal{A}}(\mu_{j}-1)+\sum\limits_{j\in\mathcal{C}}\mu_{j},$
which must be handled by non-systematic recovery sets, each of cardinality $k$. The number of non-systematic nodes in the system is $n-i$. Each non-systematic recovery set contains at least two systematic nodes since $k-(n-i)\geq 2$. Therefore, for every portion $\delta$ of $\sum\limits_{j\in\mathcal{A}}(\mu_{j}-1)+\sum\limits_{j\in\mathcal{C}}\mu_{j}$ served by non-systematic recovery sets, at least $2\delta$ must be drawn from the systematic nodes in $\mathcal{B}$, as all nodes in $\mathcal{A}$ have already been fully utilized. In other words, to serve an amount $\delta$ in the remaining requests, at least $2\delta$ must be drawn from $\sum\limits_{j\in\mathcal{B}}(1-\mu_{j}).$
On the other hand, from the condition
$i=\sum\limits_{j=1}^{k}\lambda_{j}=\sum\limits_{j\in\mathcal{A}\cup\mathcal{B}\cup\mathcal{C}}\lambda_{j},$
it follows that
$\sum\limits_{j\in\mathcal{A}}(\lambda_{j}-1)+\sum\limits_{j\in\mathcal{C}}\lambda_{j}=\sum\limits_{j\in\mathcal{B}}(1-\lambda_{j}).$
This means that after the Greedy matching allocation, the sum of the remaining requests to be served must equal the total free capacity of the systematic nodes. However, we have previously shown that to serve an amount $\delta$ in the remaining requests, at least $2\delta$ must be drawn from $\sum\limits_{j\in\mathcal{B}}(1-\mu_{j}).$ This can happen only if $\sum\limits_{j\in\mathcal{A}}(\lambda_{j}-1)+\sum\limits_{j\in\mathcal{C}}\lambda_{j}=\sum\limits_{j\in\mathcal{B}}(1-\lambda_{j})=0$ or equivalently,
$\begin{cases}\lambda_{j}=1,\quad\forall j\in\mathcal{A},\\ \lambda_{j}=0,\quad\forall j\in\mathcal{B}\cup\mathcal{C}.\end{cases}$
This implies that $\begin{array}[]{c}\boldsymbol{\mu}=(1,\ldots,1,0,\ldots,0)=\boldsymbol{\lambda},\vspace{-11mm}\\ \underbrace{\hskip 34.14322pt}_{\text{$i$ elements}}\end{array}$ contradicting our assumption that $\boldsymbol{\mu}\neq\boldsymbol{\lambda}$.
Thus, no such vector $\boldsymbol{\mu}$ can exist, proving that $\boldsymbol{\lambda}$ is the unique vector satisfying $\sum\limits_{j=1}^{k}\lambda_{j}=i$.
Moreover, the previous capacity argument also shows that $\boldsymbol{\lambda}$ can not be served by any scheme other than Greedy matching. In other words, there is only one matching w in the matching polytope such that $\boldsymbol{\lambda}=\boldsymbol{\lambda}(\textbf{w})$. ∎
VII Conclusion
In this work, we presented a rigorous analysis of the service rate region (SRR) for distributed storage systems employing MDS codes. We used graph-theoretic methods to characterize achievable rates comprehensively. By constructing a family of MDS generator matrices that vary in the number of systematic columns, we showed how increasing the number of systematic columns in the generator matrix of the same code strictly enlarges the SRR. We introduced two bounding simplices, the Maximal Matching and Maximal Achievable simplices, which provide clear geometric boundaries on feasible request rates. A key technical contribution was the proposal of a Greedy Matching allocation strategy, whose optimality we proved in several scenarios, demonstrating that certain extreme points of the SRR boundary are only attained through Greedy allocation. Using this scheme, we developed explicit characterizations of MDS-coded SRRs under various configurations of $n$ (servers), $k$ (data objects), and $i$ (systematic servers).
Looking ahead, one could extend the analysis to more general coding schemes or construct codes specifically tailored to achieve coverage for a target SRR. Future directions also include examining novel recovery constraints (e.g., locally recoverable codes) and exploiting deeper structural properties of generator matrices to refine the SRR description. By providing explicit characterizations of the SRR across diverse regimes, we anticipate that these methods will guide both practical code design and broader theoretical explorations into the interplay between code parameters and data-access performance.
Acknowledgment
The authors thank V. Lalitha for her valuable comments on an earlier draft of this work.